# Stellar properties indicating the presence of hyperons in neutron stars

### v1

@@ I Supplemental Material
EoS $M_{\rm max}$ [$M_{\odot}$] $R_{1.6}$ $\rm{[km]}$ $\Lambda_{1.6}$ Hyperons Reference
@@
ABHT(QMC-RMF1) 1.95 11.66 118.4 No [84, 110, 139]
ABHT(QMC-RMF2) 2.04 11.90 142.8 No [84, 110, 139]
ABHT(QMC-RMF3) 2.15 12.15 159.6 No [84, 110, 139]
ABHT(QMC-RMF4) 2.21 12.35 201.6 No [84, 110, 139]
@@
BL(chiral) w. un. crust 2.08 12.10 147.1 No [75, 94]
CMGO(GDFM-I) 2.30 13.89 431.1 No [138]
CMGO(GDFM-I) 2.31 12.80 233.8 No [138]
@@
GPPVA(FSU2) 2.07 13.71 346.8 No [62, 110, 118]
GPPVA(FSU2H) 2.38 13.38 354.7 No [62, 110, 118]
GPPVA(FSU2R) 2.05 12.91 260.5 No [62, 110, 118]
@@
eos1 2.21 13.08 341.9 Yes [138]
eos2 1.94 14.15 564.8 Yes [138]
eos3 2.06 14.12 479.2 Yes [138]
eos4 2.07 13.45 363.9 Yes [138]
eos5 2.07 13.53 399.8 Yes [138]
eos6 1.92 13.43 334.7 Yes [138]
eos7 2.03 13.10 295.9 Yes [138]
eos9 2.00 12.65 222.2 Yes [138]
eos0 2.18 13.19 342.4 Yes [138]
eos1 2.08 13.07 339.5 Yes [138]
eos2 2.35 14.15 564.8 Yes [138]
eos3 2.07 14.12 478.3 Yes [138]
eos4 2.08 13.45 363.4 Yes [138]
eos5 2.08 13.52 398.1 Yes [138]
eos6 1.97 13.45 340.4 Yes [138]
eos8 2.16 13.06 304.1 Yes [138]
eos0 2.56 13.19 342.4 No [138]
eos7 2.29 13.10 295.8 No [138]
(GDFM)eos1 2.36 12.89 292.7 No [138]
(GDFM)eos3 2.30 12.63 271.9 No [138]
(GDFM)eos6 2.26 12.75 240.3 No [138]
(GDFM)eos8 2.51 14.38 549.8 No [138]
(GDFM)eos11 2.92 14.18 510.1 No [138]
(GDFM)eos12 2.45 13.70 372.3 No [138]
(GDFM)eos15 2.72 14.17 446.5 No [138]
(GDFM)eos19 2.27 12.74 238.4 No [138]
(GDFM)eos20 2.65 13.63 384.4 No [138]
(GDFM)eos22 2.74 14.94 629.6 No [138]
(GDFM)eos24 2.51 14.15 401.3 No [138]
(GDFM)eos28 2.18 12.69 214.1 No [138]
(GDFM)eos30 2.87 15.16 624.8 No [138]
(GDFM)eos36 2.46 13.57 351.3 No [138]
(GDFM)eos39 2.51 14.65 583.8 No [138]
(GDFM)eos41 2.52 13.76 428.5 No [138]
(GDFM)eos43 2.77 13.86 449.0 No [138]
(GDFM)eos45 2.77 13.74 414.8 No [138]
(GDFM)eos48 2.19 13.11 264.9 No [138]
(GDFM)eos50 2.18 12.87 201.1 No [138]
(GDFM)eos52 2.08 11.66 135.8 No [138]
(GDFM)eos58 2.65 13.87 424.6 No [138]
(GDFM)eos60 2.63 14.06 456.8 No [138]
(GDFM)eos61 2.47 14.30 448.7 No [138]
(GDFM)eos64 2.18 12.99 282.7 No [138]
(GDFM)eos66 2.74 14.87 528.8 No [138]
(GDFM)eos67 2.73 14.28 552.6 No [138]
(GDFM)eos72 2.88 15.18 642.1 No [138]
(GDFM)eos78 2.88 14.53 621.8 No [138]
(GDFM)eos80 2.49 13.82 391.5 No [138]
(GDFM)eos82 2.77 14.12 540.5 No [138]
(GDFM)eos84 2.83 15.03 625.8 No [138]
(GDFM)eos85 2.63 15.02 593.5 No [138]
(GDFM)eos89 2.19 13.44 301.4 No [138]
(GDFM)eos92 2.77 14.61 584.4 No [138]
(GDFM)eos94 2.24 12.51 241.1 No [138]
(GDFM)eos10 2.55 14.09 479.2 No [138]
(GDFM)eos17 2.74 13.41 346.0 No [138]
(GDFM)eos54 2.44 12.88 304.4 No [138]
(GDFM)eos73 2.37 14.05 455.6 No [138]
(GDFM)eos99 2.33 12.32 233.0 No [138]
eos101 2.59 13.72 422.5 No [138]
eos102 2.47 12.77 301.0 No [138]
eos103 2.53 13.68 420.6 No [138]
eos105 2.59 13.16 327.7 No [138]
eos106 2.62 13.40 381.4 No [138]
eos108 2.59 13.29 366.8 No [138]
eos109 2.75 13.85 454.9 No [138]
eos110 2.39 13.02 277.4 No [138]
eos111 2.32 13.31 355.5 No [138]
eos112 2.64 13.16 311.6 No [138]
eos114 2.38 13.21 311.5 No [138]
eos117 2.35 12.65 254.3 No [138]
eos118 2.50 13.58 382.3 No [138]
eos119 2.40 13.30 335.8 No [138]
eos122 2.61 14.28 566.2 No [138]
eos125 2.41 12.62 289.9 No [138]
eos126 2.74 14.32 551.9 No [138]
eos128 2.54 13.58 423.9 No [138]
eos13 2.57 14.05 532.0 No [138]
eos130 2.69 13.65 400.9 No [138]
eos131 2.60 14.27 571.5 No [138]
eos132 2.42 12.66 248.3 No [138]
eos133 2.40 13.33 349.4 No [138]
eos135 2.68 13.91 441.6 No [138]
eos138 2.88 13.80 506.3 No [138]
eos14 2.64 13.89 483.4 No [138]
eos144 2.32 12.59 250.0 No [138]
eos147 2.51 12.76 266.5 No [138]
eos149 2.42 12.70 275.2 No [138]
eos158 2.53 14.06 489.1 No [138]
eos16 2.60 13.85 471.4 No [138]
eos160 2.72 13.27 374.9 No [138]
eos161 2.22 12.74 243.9 No [138]
eos162 2.43 12.94 317.0 No [138]
eos166 2.47 13.57 403.9 No [138]
eos168 2.53 14.13 499.3 No [138]
eos171 2.41 12.61 266.9 No [138]
eos173 2.64 13.11 379.5 No [138]
eos175 2.21 12.25 198.8 No [138]
eos176 2.47 12.86 311.1 No [138]
eos178 2.30 12.69 251.1 No [138]
eos179 2.59 13.28 370.2 No [138]
eos18 2.77 13.33 421.0 No [138]
eos182 2.79 13.41 432.4 No [138]
eos186 2.68 13.19 367.0 No [138]
eos188 2.47 13.26 316.2 No [138]
eos191 2.46 14.27 536.7 No [138]
eos193 2.52 13.46 363.5 No [138]
eos196 2.48 12.75 303.7 No [138]
eos198 2.48 12.79 297.0 No [138]
eos2 2.67 14.15 564.8 No [138]
eos200 2.55 13.68 395.5 No [138]
eos203 2.31 13.39 380.6 No [138]
eos207 2.52 14.08 504.9 No [138]
eos208 2.59 13.16 354.5 No [138]
eos211 2.43 13.05 331.3 No [138]
eos213 2.59 14.12 510.3 No [138]
eos216 2.75 13.46 404.7 No [138]
eos223 2.42 13.51 351.3 No [138]
eos224 2.63 14.06 496.5 No [138]
eos23 2.46 13.58 371.6 No [138]
eos231 2.46 13.50 373.9 No [138]
eos232 2.65 13.05 322.3 No [138]
eos234 2.52 12.72 296.1 No [138]
eos26 2.30 13.38 341.7 No [138]
eos27 2.67 13.36 416.8 No [138]
eos29 2.78 13.49 397.3 No [138]
eos33 2.18 13.15 294.6 No [138]
eos35 2.35 13.68 387.3 No [138]
eos37 2.40 13.73 411.3 No [138]
eos4 2.42 13.46 365.1 No [138]
eos42 2.41 12.61 300.6 No [138]
eos46 2.29 13.25 348.6 No [138]
eos47 2.37 13.08 308.5 No [138]
eos49 2.48 14.08 513.5 No [138]
eos5 2.47 13.53 400.4 No [138]
eos53 2.43 14.17 492.0 No [138]
eos54 2.37 12.73 288.5 No [138]
eos55 2.76 13.83 431.2 No [138]
eos56 2.46 12.53 268.3 No [138]
eos57 2.50 13.36 348.4 No [138]
eos59 2.69 13.31 362.4 No [138]
eos63 2.38 13.65 376.7 No [138]
eos65 2.73 13.20 377.1 No [138]
eos68 2.24 12.55 227.3 No [138]
eos69 2.76 14.18 566.6 No [138]
eos71 2.43 12.93 310.7 No [138]
eos76 2.32 12.67 258.0 No [138]
eos77 2.37 12.70 273.8 No [138]
eos78 2.85 13.42 435.4 No [138]
eos79 2.29 12.87 248.7 No [138]
eos81 2.86 13.88 472.0 No [138]
eos83 2.99 14.22 595.6 No [138]
eos87 2.73 14.36 577.6 No [138]
eos88 2.68 13.27 385.3 No [138]
eos9 2.22 12.68 227.5 No [138]
eos91 2.42 13.38 383.4 No [138]
eos93 2.54 12.97 345.4 No [138]
eos97 2.86 14.01 512.0 No [138]
eos98 2.73 14.11 545.6 No [138]

### v2

@@ Appendix A EoS models and additional plots
EoS $_{\rm max} [MM_{\odot}$] $R_{1.6}$ $\rm{[km]}$ $\Lambda_{1.6}$ Hyperons Reference
@@
ABHT(QMC-RMF1) 1.95 11.66 118.4 No [92, 118, 147, 148]
ABHT(QMC-RMF2) 2.04 11.90 142.8 No [92, 118, 147, 148]
ABHT(QMC-RMF3) 2.15 12.15 159.6 No [92, 118, 147, 148]
ABHT(QMC-RMF4) 2.21 12.35 201.6 No [92, 118, 147, 148]
@@
BL(chiral) w. un. crust 2.08 12.10 147.1 No [82, 149, 158]
CMGO(GDFM-II) 2.30 13.89 431.1 No [146, 149]
CMGO(GDFM-I) 2.31 12.80 233.8 No [146, 149]
@@
GPPVA(FSU2) 2.07 13.71 346.8 No [150, 118, 126, 151]
GPPVA(FSU2H) 2.38 13.38 354.7 No [150, 118, 126, 29]
GPPVA(FSU2R) 2.05 12.91 260.5 No [150, 118, 126, 29]
@@
eos1 2.21 13.08 341.9 Yes [157, 158]
eos2 1.94 14.15 564.8 Yes [157, 158]
eos3 2.06 14.12 479.2 Yes [157, 158]
eos4 2.07 13.45 363.9 Yes [157, 158]
eos5 2.07 13.53 399.8 Yes [157, 158]
eos6 1.92 13.43 334.7 Yes [157, 158]
eos7 2.03 13.10 295.9 Yes [157, 158]
eos9 2.00 12.65 222.2 Yes [157, 158]
eos0 2.18 13.19 342.4 Yes [157, 158]
eos1 2.08 13.07 339.5 Yes [157, 158]
eos2 2.35 14.15 564.8 Yes [157, 158]
eos3 2.07 14.12 478.3 Yes [157, 158]
eos4 2.08 13.45 363.4 Yes [157, 158]
eos5 2.08 13.52 398.1 Yes [157, 158]
eos6 1.97 13.45 340.4 Yes [157, 158]
eos8 2.16 13.06 304.1 Yes [157, 158]
eos0 2.56 13.19 342.4 No [146, 158]
eos7 2.29 13.10 295.8 No [146, 158]
(GDFM)eos1 2.36 12.89 292.7 No [146, 158]
(GDFM)eos3 2.30 12.63 271.9 No [146, 158]
(GDFM)eos6 2.26 12.75 240.3 No [146, 158]
(GDFM)eos8 2.51 14.38 549.8 No [146, 158]
(GDFM)eos11 2.92 14.18 510.1 No [146, 158]
(GDFM)eos12 2.45 13.70 372.3 No [146, 158]
(GDFM)eos15 2.72 14.17 446.5 No [146, 158]
(GDFM)eos19 2.27 12.74 238.4 No [146, 158]
(GDFM)eos20 2.65 13.63 384.4 No [146, 158]
(GDFM)eos22 2.74 14.94 629.6 No [146, 158]
(GDFM)eos24 2.51 14.15 401.3 No [146, 158]
(GDFM)eos28 2.18 12.69 214.1 No [146, 158]
(GDFM)eos30 2.87 15.16 624.8 No [146, 158]
(GDFM)eos36 2.46 13.57 351.3 No [146, 158]
(GDFM)eos39 2.51 14.65 583.8 No [146, 158]
(GDFM)eos41 2.52 13.76 428.5 No [146, 158]
(GDFM)eos43 2.77 13.86 449.0 No [146, 158]
(GDFM)eos45 2.77 13.74 414.8 No [146, 158]
(GDFM)eos48 2.19 13.11 264.9 No [146, 158]
(GDFM)eos50 2.18 12.87 201.1 No [146, 158]
(GDFM)eos52 2.08 11.66 135.8 No [146, 158]
(GDFM)eos58 2.65 13.87 424.6 No [146, 158]
(GDFM)eos60 2.63 14.06 456.8 No [146, 158]
(GDFM)eos61 2.47 14.30 448.7 No [146, 158]
(GDFM)eos64 2.18 12.99 282.7 No [146, 158]
(GDFM)eos66 2.74 14.87 528.8 No [146, 158]
(GDFM)eos67 2.73 14.28 552.6 No [146, 158]
(GDFM)eos72 2.88 15.18 642.1 No [146, 158]
(GDFM)eos78 2.88 14.53 621.8 No [146, 158]
(GDFM)eos80 2.49 13.82 391.5 No [146, 158]
(GDFM)eos82 2.77 14.12 540.5 No [146, 158]
(GDFM)eos84 2.83 15.03 625.8 No [146, 158]
(GDFM)eos85 2.63 15.02 593.5 No [146, 158]
(GDFM)eos89 2.19 13.44 301.4 No [146, 158]
(GDFM)eos92 2.77 14.61 584.4 No [146, 158]
(GDFM)eos94 2.24 12.51 241.1 No [146, 158]
(GDFM)eos10 2.55 14.09 479.2 No [146, 158]
(GDFM)eos17 2.74 13.41 346.0 No [146, 158]
(GDFM)eos54 2.44 12.88 304.4 No [146, 158]
(GDFM)eos73 2.37 14.05 455.6 No [146, 158]
(GDFM)eos99 2.33 12.32 233.0 No [146, 158]
eos101 2.59 13.72 422.5 No [146, 158]
eos102 2.47 12.77 301.0 No [146, 158]
eos103 2.53 13.68 420.6 No [146, 158]
eos105 2.59 13.16 327.7 No [146, 158]
eos106 2.62 13.40 381.4 No [146, 158]
eos108 2.59 13.29 366.8 No [146, 158]
eos109 2.75 13.85 454.9 No [146, 158]
eos110 2.39 13.02 277.4 No [146, 158]
eos111 2.32 13.31 355.5 No [146, 158]
eos112 2.64 13.16 311.6 No [146, 158]
eos114 2.38 13.21 311.5 No [146, 158]
eos117 2.35 12.65 254.3 No [146, 158]
eos118 2.50 13.58 382.3 No [146, 158]
eos119 2.40 13.30 335.8 No [146, 158]
eos122 2.61 14.28 566.2 No [146, 158]
eos125 2.41 12.62 289.9 No [146, 158]
eos126 2.74 14.32 551.9 No [146, 158]
eos128 2.54 13.58 423.9 No [146, 158]
eos13 2.57 14.05 532.0 No [146, 158]
eos130 2.69 13.65 400.9 No [146, 158]
eos131 2.60 14.27 571.5 No [146, 158]
eos132 2.42 12.66 248.3 No [146, 158]
eos133 2.40 13.33 349.4 No [146, 158]
eos135 2.68 13.91 441.6 No [146, 158]
eos138 2.88 13.80 506.3 No [146, 158]
eos14 2.64 13.89 483.4 No [146, 158]
eos144 2.32 12.59 250.0 No [146, 158]
eos147 2.51 12.76 266.5 No [146, 158]
eos149 2.42 12.70 275.2 No [146, 158]
eos158 2.53 14.06 489.1 No [146, 158]
eos16 2.60 13.85 471.4 No [146, 158]
eos160 2.72 13.27 374.9 No [146, 158]
eos161 2.22 12.74 243.9 No [146, 158]
eos162 2.43 12.94 317.0 No [146, 158]
eos166 2.47 13.57 403.9 No [146, 158]
eos168 2.53 14.13 499.3 No [146, 158]
eos171 2.41 12.61 266.9 No [146, 158]
eos173 2.64 13.11 379.5 No [146, 158]
eos175 2.21 12.25 198.8 No [146, 158]
eos176 2.47 12.86 311.1 No [146, 158]
eos178 2.30 12.69 251.1 No [146, 158]
eos179 2.59 13.28 370.2 No [146, 158]
eos18 2.77 13.33 421.0 No [146, 158]
eos182 2.79 13.41 432.4 No [146, 158]
eos186 2.68 13.19 367.0 No [146, 158]
eos188 2.47 13.26 316.2 No [146, 158]
eos191 2.46 14.27 536.7 No [146, 158]
eos193 2.52 13.46 363.5 No [146, 158]
eos196 2.48 12.75 303.7 No [146, 158]
eos198 2.48 12.79 297.0 No [146, 158]
eos2 2.67 14.15 564.8 No [146, 158]
eos200 2.55 13.68 395.5 No [146, 158]
eos203 2.31 13.39 380.6 No [146, 158]
eos207 2.52 14.08 504.9 No [146, 158]
eos208 2.59 13.16 354.5 No [146, 158]
eos211 2.43 13.05 331.3 No [146, 158]
eos213 2.59 14.12 510.3 No [146, 158]
eos216 2.75 13.46 404.7 No [146, 158]
eos223 2.42 13.51 351.3 No [146, 158]
eos224 2.63 14.06 496.5 No [146, 158]
eos23 2.46 13.58 371.6 No [146, 158]
eos231 2.46 13.50 373.9 No [146, 158]
eos232 2.65 13.05 322.3 No [146, 158]
eos234 2.52 12.72 296.1 No [146, 158]
eos26 2.30 13.38 341.7 No [146, 158]
eos27 2.67 13.36 416.8 No [146, 158]
eos29 2.78 13.49 397.3 No [146, 158]
eos33 2.18 13.15 294.6 No [146, 158]
eos35 2.35 13.68 387.3 No [146, 158]
eos37 2.40 13.73 411.3 No [146, 158]
eos4 2.42 13.46 365.1 No [146, 158]
eos42 2.41 12.61 300.6 No [146, 158]
eos46 2.29 13.25 348.6 No [146, 158]
eos47 2.37 13.08 308.5 No [146, 158]
eos49 2.48 14.08 513.5 No [146, 158]
eos5 2.47 13.53 400.4 No [146, 158]
eos53 2.43 14.17 492.0 No [146, 158]
eos54 2.37 12.73 288.5 No [146, 158]
eos55 2.76 13.83 431.2 No [146, 158]
eos56 2.46 12.53 268.3 No [146, 158]
eos57 2.50 13.36 348.4 No [146, 158]
eos59 2.69 13.31 362.4 No [146, 158]
eos63 2.38 13.65 376.7 No [146, 158]
eos65 2.73 13.20 377.1 No [146, 158]
eos68 2.24 12.55 227.3 No [146, 158]
eos69 2.76 14.18 566.6 No [146, 158]
eos71 2.43 12.93 310.7 No [146, 158]
eos76 2.32 12.67 258.0 No [146, 158]
eos77 2.37 12.70 273.8 No [146, 158]
eos78 2.85 13.42 435.4 No [146, 158]
eos79 2.29 12.87 248.7 No [146, 158]
eos81 2.86 13.88 472.0 No [146, 158]
eos83 2.99 14.22 595.6 No [146, 158]
eos87 2.73 14.36 577.6 No [146, 158]
eos88 2.68 13.27 385.3 No [146, 158]
eos9 2.22 12.68 227.5 No [146, 158]
eos91 2.42 13.38 383.4 No [146, 158]
eos93 2.54 12.97 345.4 No [146, 158]
eos97 2.86 14.01 512.0 No [146, 158]
eos98 2.73 14.11 545.6 No [146, 158]
@@
Figure 8 shows $\frac{d^{2}R}{dM^{2}}(M)$ and $\frac{d^{2}\Lambda}{dM^{2}}(M)$ at a fixed reference mass $M=1.6\,M_{\odot}$ similar to Figs. 3 and 5 in the main paper. For the right panel we adopt $\Delta=0.15~M_{\odot}$ because the discretization errors become comparable to errors stemming from the finite precision in $\Lambda$ if $\Delta=0.25~M_{\odot}$ is used.
@@
Figure 8: Second derivative of $R(M)$ (left) and $\Lambda(M)$ (right) with respect to mass $M$ at a fixed reference mass $M=1.6\,M_{\odot}$ as function of $R(M=1.6\,M_{\odot})$ (left) and $\Lambda(M=1.6\,M_{\odot})$, respectively. Symbols have the same meaning as in Fig. 3 of the main paper. We adopt $\delta R=100$ m and $\Delta M=0.25\,M_{\odot}$ for the left panel and $\delta\Lambda=100/1.6^{5}$ and $\Delta M=0.15\,M_{\odot}$ for the right panel. See main text for more details.